# Evaluating many-body stabilizer Rényi entropy by sampling reduced Pauli strings: singularities, volume law, and nonlocal magic

### v2

@@ Appendix A Nonlocal update and autocorrelation time for simulating $Q$ in the TFI model
$H_{-1,j}=h(\sigma_{j}^{+}+\sigma_{j}^{-})$ (33)
$H_{0,j}=hI_{j}$ (34)
$H_{1,\overline{jk}}=J(Z_{j}Z_{k}+I_{j}I_{k})$ (35)
@@
$\tau_{n_{J}}^{\text{int}}:=\frac{1}{2}+\sum_{t=1}^{\infty}A_{n_{J}}(t)$ (36)
@@
is the autocorrelation function with $u$ and $t$ to denote different MC times [88, 73]. $A_{n_{J}}(t)$ is a convex function and will decay exponetially when $t$ is large, and the sum in Eq. (36) is truncated when $A_{n_{J}}(t)$ becomes negligibly small in practice [88]. The results are shown in Fig. 11, which indicate the efficiency of this modifed algorithm for nonlocal updates.
@@ Appendix B Estimators in reweight-annealing
$\begin{split}\frac{Q(\lambda_{k-1})}{Q(\lambda_{k})}=&\frac{\sum_{c}W_{c}(\lambda_{k-1})}{\sum_{c}W_{c}(\lambda_{k})}\\ =&\frac{\sum_{c}\frac{W_{c}(\lambda_{k-1})}{W_{c}(\lambda_{k})}W_{c}(\lambda_{k})}{\sum_{c}W_{c}(\lambda_{k})}\\ =&\bigg{\langle}\frac{W(\lambda_{k-1})}{W(\lambda_{k})}\bigg{\rangle}_{\lambda_{k},Q}\end{split}$ (38)
which averages the weight ratios for the two parameters $\lambda_{k-1}$ and $\lambda_{k}$. Eq. (38) is called the reweighting trick [89, 38]. If $\lambda\equiv\beta$, then according to Eq. (13) and Eq. (14), the ratio $W(\beta_{k-1})/W(\beta_{k})$ in Eq. (38) becomes $(\beta_{k-1}/\beta_{k})^{n_{\text{tot}}}$, where $n_{\text{tot}}$ is the total number of operators in the four replicas. Simialrly, we can obtain the reweighting ratio when $\lambda\equiv J$.

### v3

@@ Appendix A Nonlocal update and autocorrelation time for simulating $Q$ in the TFI model
$H_{-1,j}=h(\sigma_{j}^{+}+\sigma_{j}^{-}),$ (33)
$H_{0,j}=hI_{j},$ (34)
$H_{1,\overline{jk}}=J(Z_{j}Z_{k}+I_{j}I_{k}),$ (35)
@@
$\tau_{n_{J}}^{\text{int}}:=\frac{1}{2}+\sum_{t=1}^{\infty}A_{n_{J}}(t),$ (36)
@@
is the autocorrelation function with $u$ and $t$ to denote different MC times [100, 80]. $A_{n_{J}}(t)$ is a convex function and will decay exponentially when $t$ is large, and the sum in Eq. (36) is truncated when $A_{n_{J}}(t)$ becomes negligibly small in practice [100]. The results are shown in Fig. 12, which indicate the efficiency of this modified algorithm for nonlocal updates.
@@ Appendix B Estimators in reweight-annealing
$\begin{split}\frac{Q(\lambda_{k-1})}{Q(\lambda_{k})}=&\frac{\sum_{c}W_{c}(\lambda_{k-1})}{\sum_{c}W_{c}(\lambda_{k})}\\ =&\frac{\sum_{c}\frac{W_{c}(\lambda_{k-1})}{W_{c}(\lambda_{k})}W_{c}(\lambda_{k})}{\sum_{c}W_{c}(\lambda_{k})}\\ =&\bigg{\langle}\frac{W(\lambda_{k-1})}{W(\lambda_{k})}\bigg{\rangle}_{\lambda_{k},Q},\end{split}$ (38)
which averages the weight ratios for the two parameters $\lambda_{k-1}$ and $\lambda_{k}$. Eq. (38) is called the reweighting trick [101, 38]. If $\lambda\equiv\beta$, then according to Eq. (13) and Eq. (14), the ratio $W(\beta_{k-1})/W(\beta_{k})$ in Eq. (38) becomes $(\beta_{k-1}/\beta_{k})^{n_{\text{tot}}}$, where $n_{\text{tot}}$ is the total number of operators in the four replicas. Similarly, we can obtain the reweighting ratio when $\lambda\equiv J$.